OpenStudy (anonymous):

step by step pls 1/4s = 8

OpenStudy (anonymous):

solve for s?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/4s=8 (1/4)*(4)s=8*4 1*s=32 s=32

OpenStudy (anonymous):

1/(4s)=8 1=8*4s=32s, so s=1/32

OpenStudy (saifoo.khan):

Please write the correct postition of S.

OpenStudy (saifoo.khan):

Position*

OpenStudy (anonymous):

is the s a part of the denominator or not?

OpenStudy (saifoo.khan):

Yeah.

OpenStudy (anonymous):

is it (1/4)s=8 or 1/(4s)=8?